Southwest Dry Cleaners believes that it will need equipment in 10 yeas. The equipment will cost 26,000. What lump sum should be invested today at 6% compounded semiannually to yield $26,000?


1
Expert's answer
2021-07-15T11:25:37-0400

"A=P(1+\frac {r} {n}) ^{nt}"


"P=\frac {A} {(1+\frac {r} {n}) ^{nt} }"


"P=\frac {26000}{(1+\frac {0.06} {2}) ^{(10×2)}}"


"P=\frac {26000}{(1+0.03)^{20}}"


"P=\$14,395.57"